OpenStudy (anonymous):

please solve for m (m+8/m^2-3m-4) - (8/m^2-8m+16) = (m-8)/m^2-3m-4

OpenStudy (anonymous):

first thing factorize all denominators

OpenStudy (anonymous):

how can i factorize the denominator because for example m^2-3m-4.. nothing will go into all three of these terms

OpenStudy (anonymous):

rewrite -3m as -4m+m and factorize by grouping

OpenStudy (anonymous):

ok,i'll do this one for you: m^2-3m-4: you look for 2 numbers whose product is -4 and whose sum is -3:-4,1 now rewrite it as m^2-4m+m-4 now factor each two terms alone,that's grouping m(m-4)+(m-4)=(m-4)(m+1)

OpenStudy (anonymous):

like a quadratic equation?

OpenStudy (anonymous):

yes,there are many methods to factorize these,i don't know which one you use

OpenStudy (anonymous):

i'm not really sure what to do with this problem at that point. should i divide the (m-4)(m+1) by the other side... can i even do this since it is already a fraction??

OpenStudy (anonymous):

just factorize the other denominators for now

OpenStudy (anonymous):

can you do it?

OpenStudy (anonymous):

yes i'm working on factorizing right now

OpenStudy (anonymous):

ok,take your time:)

OpenStudy (anonymous):

m(m-4)+(m-4)=(m-4)(m+1) m(m-4)-4(m-4)=(m-4)(m-4) m(m-4)+(m+4)=(m-4)(m+1)

OpenStudy (anonymous):

the last one should be m(m-4)+(m-4)=(m-4)(m+1)

OpenStudy (anonymous):

that's the one i did! factorize m^2-8m+16 and m^2-3m-4

OpenStudy (anonymous):

sorry i can't believe i did the same one three times. all of this algebra is getting to my head. okay i'm doing it now

OpenStudy (anonymous):

relax..

OpenStudy (anonymous):

so for the m^2-8m+16 -4*-4 give me 16... and -4+-4 = 8.... m^2-4m-4m+16.. m(m-4)-4(m-4)=(m-4)(m-4)

OpenStudy (anonymous):

m^2-3m-4 -4*1 = -4 and added gives me -3 so.. m^2-4m+1m-4 m(m-4)+(m-4)=(m-4)(m+1)

OpenStudy (anonymous):

good